Assuming I am given a number 12. Find out by calculations if Four more than half the number is eight more than twenty-four times the reciprocal of the number.
Answer by algebrahouse.com(1659)

0.5(12) + 4 = 8 + 24(1/12) {four more than half 12 is 8 more than 24 times reciprocal}
6 + 4 = 8 + 2 {multiplied}
10 = 10 {added}

true statement...answer to question is yes

For more help from me, visit: www.algebrahouse.com